OpenStudy (anonymous):

Differentiate the function y = f(x) = x^3 arctan(x).

OpenStudy (anonymous):

product rule: \(\large [x^3 \cdot arctanx]'=[x^3]'\cdot arctanx + x^3 \cdot [arctanx]' \)

OpenStudy (anonymous):

thanks!